wikki4deal:

Q: Find the product with the exponent in simplest form. Then, identify the values of x and y. 6^1/3x6^1/4=6^x/y X= ,y=

ch3wyc00ki3:

so 6 to the first power is 6

ch3wyc00ki3:

and then do 3*6=18

ch3wyc00ki3:

then divide 6/18

ch3wyc00ki3:

do u know how to do the rest

jhonyy9:

use this formula x^a * x^b = x^(a+b)